# Scalable and Cost-Efficient de Novo Template-Based Molecular Generation

### v1

@@ Appendix E Extended Synthesis Cost Guidance Analysis
In this section, we extend the analysis from Section 4.3 with results in the SMALL setting. Figure 8 shows that similarly to the MEDIUM setting, Synthesis Cost Guidance reduces the average length of the trajectories and the average price of the fragments used also in the SMALL setting. Interestingly, it does not significantly limit the usage of fragments, which is likely due to their shortage and relatively uniform cost. We also see that the reaction yield does not play a role in reducing the cost.
@@
Appendix F Further Comparison to Baselines
In this section, we extend our comparison of SCENT to established baseline models from the literature: REINVENT [20] and Graph GA [53]. Figure 9 shows a comparative analysis of the molecular properties generated by each method, filtered through AIZynthFinder [48] to retain only retrosynthetically accessible compounds. SCENT achieves a reward comparable to REINVENT, while outperforming it in terms of diversity and novelty. Graph GA generates highly diverse molecules, but these are often challenging to synthesize, as indicated by its low support in the final sub-plot. Importantly, both SCENT and REINVENT lie on the Pareto front of the reward–novelty trade-off: SCENT is biased toward higher novelty, whereas REINVENT tends toward higher reward. LABEL:tab:synth_metrics_1 and LABEL:tab:synth_metrics_2 further support this comparison, confirming that SCENT generates diverse, high-quality, and synthetically accessible molecules.
@@ Appendix F Further Comparison to Baselines
Template-based RGFN 7.34 $\pm$0.06 0.77 $\pm$0.02 0.58 $\pm$0.01 0.647 $\pm$0.037
@@
Appendix G Ablations
G.1 Decomposability Guidance Ablation
Decomposability Guidance, described in Section 3.2.2, is an integral part of our SCENT model. Leveraging a machine learning model, it predicts which molecules can be further decomposed in the backward transition (retrosynthesized) and discards the transitions that would otherwise lead to invalid states. Table 7 shows that this component is crucial for the performance of SCENT.
@@ G.1 Decomposability Guidance Ablation
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
@@
G.2 Synthesibility Cost Guidance Ablations
G.2.1 Cost Prediction Model Ablation
In this section, we demonstrate the importance of accurately estimating the true cost $c$ using the machine learning model $c^{\prime}$ for the optimal performance of SCENT. We compare our model with a low-quality cost estimator that outputs a constant value for every input molecule, where this constant is the mean of all the building blocks in the given scenario. It is important to note that we only modify the recursive component of the $C^{S}$ function from Equation 5, keeping the fragment costs unchanged. As shown in Table 8, the use of a constant value can reduce costs and improve SCENT’s performance in MEDIUM and LARGE settings. However, machine learning-based cost estimation significantly outperforms this approach, providing a substantial performance boost, particularly in cost reduction. From this, we conclude that the quality of the cost prediction model used in Cost Guidance is critical to the performance of SCENT.
@@ G.2.1 Cost Prediction Model Ablation
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
@@
LARGE w/o Cost Guidance 7171 $\pm$291 8767 $\pm$429 7.13 $\pm$0.12 1678 $\pm$63
@@
G.2.2 Comparison of Cost-Reducing Methods
In this section, we compare our cost-guided SCENT with other methods to reduce the synthesis cost. The first method is to use the forward synthesis cost as a reward function component (we denote this approach as the "cost reward"). The second method that we tested is trained on subsets of the building block library $M$ with the top $n\%$ of the most expensive molecules being removed. While both approaches can effectively reduce the cost of proposed molecules, they lag behind cost-guided SCENT in terms of online discovery metrics, as can be seen in Figure 10. This suggests that cost biasing strikes the perfect balance between cost-effectiveness and diversity as it is able to consistently find more modes.
@@ G.2.2 Comparison of Cost-Reducing Methods
G.3 Dynamic Library Ablation
Figure 11 illustrates how the performance of the reward-based Dynamic Library scales with the size of the initial set of building blocks. The improvement over the baseline (SCENT (C+P)) is substantial when the initial library is small but diminishes as its size increases. We also conduct an ablation study on the library-building mechanism, comparing our method—selecting the top $L$ molecules from the reward-sorted buffer—to a variant where $L$ molecules are randomly sampled from the buffer. Our results indicate that the selection mechanism plays a crucial role and is the primary driver of performance improvements.
@@ G.3 Dynamic Library Ablation
G.4 Exploitation Penalty vs Other Exploration Techniques
In this section, we compare our Exploitation Penalty to the widely used $\epsilon$-greedy exploration where $\epsilon=0.05$. Additionally, we "decay" variants that linearly decrease the exploration parameters ($\epsilon$ for $\epsilon$-greedy, and $\gamma$ for exploitation penalty) so that they reach 00 at the $1000$-th iteration. Figure 12 compares the exploration techniques on three variations of SCENT: with/without Cost Guidance, and with Dynamic Library. We observe that exploitation penalty with linear decaying (penalty decaying) helps our cost guided SCENT the most, successfully mitigating its exploitative nature. Dynamic Library also benefits from using an exploitation penalty, possibly because the penalty explicitly encourages the model to visit newly added molecules. Table 9 shows that our Exploration Penalty helps the cost-guided SCENT in exploration, especially when accompanied with Dynamic Library.
@@ G.4 Exploitation Penalty vs Other Exploration Techniques
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
w/o cost guidance disabled 613 $\pm$20 7599 $\pm$233 7.42 $\pm$0.05 39.3 $\pm$3.9
@@
cost guidance disabled 509 $\pm$24 5794 $\pm$337 7.4 $\pm$0.08 23.9 $\pm$7.6
@@
cost guidance + dyn. library disabled 555 $\pm$43 7098 $\pm$770 7.56 $\pm$0.04 20.5 $\pm$2.6
@@
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL SCENT (w/o C) 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0

### v2

@@ Appendix H Extended Synthesis Cost Guidance Analysis
In this section, we extend the analysis from Section˜4.3 with results in the SMALL setting. Figure˜12 shows that similarly to the MEDIUM setting, Synthesis Cost Guidance reduces the average length of the trajectories and the average price of the fragments used also in the SMALL setting. Interestingly, it does not significantly limit the usage of fragments, which is likely due to their shortage and relatively uniform cost. We also see that the reaction yield does not play a role in reducing the cost.
@@
Appendix I Further Comparison to Baselines
In this section, we extend our comparison of SCENT to established baseline models from the literature: REINVENT [20] and Graph GA [53]. Figure˜13 shows a comparative analysis of the molecular properties generated by each method, filtered through AIZynthFinder [48] to retain only retrosynthetically accessible compounds. SCENT achieves a reward comparable to REINVENT, while outperforming it in terms of diversity and novelty. Graph GA generates highly diverse molecules, but these are often challenging to synthesize, as indicated by its low support in the final sub-plot. Importantly, both SCENT and REINVENT lie on the Pareto front of the reward–novelty trade-off: SCENT is biased toward higher novelty, whereas REINVENT tends toward higher reward. LABEL:tab:synth_metrics_1 and LABEL:tab:synth_metrics_2 further support this comparison, confirming that SCENT generates diverse, high-quality, and synthetically accessible molecules.
@@ Appendix I Further Comparison to Baselines
Template-based
RGFN 7.34 $\pm$0.06 0.77 $\pm$0.02 0.58 $\pm$0.01 0.647 $\pm$0.037
@@
Appendix J Ablations
J.1 Decomposability Guidance Ablation
Decomposability Guidance, described in Section˜3.2.2, is an integral part of our SCENT model. Leveraging a machine learning model, it predicts which molecules can be further decomposed in the backward transition (retrosynthesized) and discards the transitions that would otherwise lead to invalid states. Table˜9 shows that this component is crucial for the performance of SCENT.
@@
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
@@
J.2 Synthesibility Cost Guidance Ablations
J.2.1 Cost Prediction Model Ablation
In this section, we demonstrate the importance of accurately estimating the true cost $c$ using the machine learning model $c^{\prime}$ for the optimal performance of SCENT. We compare our model with a low-quality cost estimator that outputs a constant value for every input molecule, where this constant is the mean of all the building blocks in the given scenario. It is important to note that we only modify the recursive component of the $C^{S}$ function from Equation˜5, keeping the fragment costs unchanged. As shown in Table˜10, the use of a constant value can reduce costs and improve SCENT’s performance in MEDIUM and LARGE settings. However, machine learning-based cost estimation significantly outperforms this approach, providing a substantial performance boost, particularly in cost reduction. From this, we conclude that the quality of the cost prediction model used in Cost Guidance is critical to the performance of SCENT.
@@
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
@@
LARGE w/o Cost Guidance $71$ $\pm$291 8767 $\pm$429 7.13 $\pm$0.12 1678 $\pm$63
@@
J.2.2 Comparison of Cost-Reducing Methods
In this section, we compare our cost-guided SCENT with other methods to reduce the synthesis cost. The first method is to use the forward synthesis cost as a reward function component (we denote this approach as the "cost reward"). The second method that we tested is trained on subsets of the building block library $M$ with the top $n\%$ of the most expensive molecules being removed. While both approaches can effectively reduce the cost of proposed molecules, they lag behind cost-guided SCENT in terms of online discovery metrics, as can be seen in Figure˜14. This suggests that cost biasing strikes the perfect balance between cost-effectiveness and diversity as it is able to consistently find more modes.
@@
J.2.3 Update of Building Block Costs
We conducted an experiment on SCENT (C+D+P) to evaluate system robustness to changes in fragment costs.
At iteration 3000, we randomly reshuffled fragment costs and continued training without resetting model parameters, only updating the costs estimates used to train the cost predictor. Plots showing the evolution of cost and reward over time can be found in Figure˜15. As expected, the trajectory cost spiked immediately after the change but quickly decreased, converging to the same level reached by a model trained from scratch with the new cost assignments (within  1000 iterations). The conclusion is that SCEN (C+D+P) does not require full retraining when costs change. Updating the cost predictor is sufficient.
Figure 15: The evolution of cost and reward over training iterations. “Standard” refers to the method where costs were shuffled at the start of training. “Shuffled” refers to the method where shuffling occurred at the 3000th iteration. Results are reported on 1500 validation trajectories and averaged across 3 random seeds.
J.3 Dynamic Library Ablation
Figure˜16 illustrates how the performance of the reward-based Dynamic Library scales with the size of the initial set of building blocks. The improvement over the baseline (SCENT (C+P)) is substantial when the initial library is small but diminishes as its size increases. We also conduct an ablation study on the library-building mechanism, comparing our method—selecting the top $L$ molecules from the reward-sorted buffer—to a variant where $L$ molecules are randomly sampled from the buffer. Our results indicate that the selection mechanism plays a crucial role and is the primary driver of performance improvements.
@@
J.4 Exploitation Penalty vs Other Exploration Techniques
In this section, we compare our Exploitation Penalty to the widely used $\epsilon$-greedy exploration where $\epsilon=0.05$. Additionally, we "decay" variants that linearly decrease the exploration parameters ($\epsilon$ for $\epsilon$-greedy, and $\gamma$ for exploitation penalty) so that they reach 0 at the $1000$-th iteration. Figure˜17 compares the exploration techniques on three variations of SCENT: with/without Cost Guidance, and with Dynamic Library. We observe that exploitation penalty with linear decaying (penalty decaying) helps our cost guided SCENT the most, successfully mitigating its exploitative nature. Dynamic Library also benefits from using an exploitation penalty, possibly because the penalty explicitly encourages the model to visit newly added molecules. Table˜11 shows that our Exploitation Penalty helps the cost-guided SCENT in exploration, especially when accompanied with Dynamic Library.
@@
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
w/o
cost guidance
disabled 613 $\pm$20 7599 $\pm$233 7.42 $\pm$0.05 39.3 $\pm$3.9
@@
cost guidance
disabled 509 $\pm$24 5794 $\pm$337 7.4 $\pm$0.08 23.9 $\pm$7.6
@@
cost guidance
+ dyn. library
disabled 555 $\pm$43 7098 $\pm$770 7.56 $\pm$0.04 20.5 $\pm$2.6
@@
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
SMALL
SCENT (w/o C) 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0